OpenStudy (anonymous):

Please help! solve for y. x+2y=5

OpenStudy (zzr0ck3r):

subtract x from both sides, then divide both sides by 2

OpenStudy (anonymous):

is it 2.5?

OpenStudy (anonymous):

y=5+x/2

OpenStudy (anonymous):

how will I be able to write this answer in this format? y=(_)/_

OpenStudy (anonymous):

see the question is solve for y therefore we have to find the value of y in terms of x

OpenStudy (anonymous):

so shift everything except y to the right!

OpenStudy (anonymous):

i dont get it.

OpenStudy (anonymous):

We need to solve for y, but we can't COMPLETELY solve for y unless we know x too. x+2y=5 ---Subtract x from both sides 2y=-x+5 ---Divide everything by 2 y = -1/2x +5/2 Does this make sense?

OpenStudy (anonymous):

yes so i just have to solve it now?

OpenStudy (anonymous):

@NinjaDevo

OpenStudy (anonymous):

Well...you have already solved for y, because we have y = something Think about it this way. When we usually solve for y, we have something like y = 6 It doesn't necessarily have to be a number though, the solution could be y = 2x-z-9+r+w It doesn't matter, as long as we have y = something.